OpenStudy (anonymous):

Which set of line segments with the given lengths can be formed into a right triangle? five, six, eleven five, nine, ten five, thirteen, eighteen five, twelve, thirteen

OpenStudy (anonymous):

you would have to try them all out using the pythagorean theorem

OpenStudy (anonymous):

and see, which one gives you a reasonable answer

OpenStudy (anonymous):

my brother and i used that but we got different answers would you try to solve it?

OpenStudy (anonymous):

25 + 36 = 121 ? 25 + 81 = 100 ? 25 + 169 = 324 ? 25 + 144 = 169 ? Yes!

OpenStudy (anonymous):

i just worked it out, and i also got d!

hero (hero):

There's a much easier way of course. You don't need to use pythagorean theorem. Pythagorean Theorem doesn't always work for this.